# Deep learning reveals key predictors of thermal conductivity in covalent organic frameworks

### v2

@@ Results
In this study, we leveraged the pre-trained transformer encoder of the PMTransformer model to perform transfer learning for the prediction of the thermal conductivity of COFs. We fine-tune the transformer encoder model using the pre-trained weights as a starting point and jointly train a prediction head based on a multi-layer perceptron architecture to map the output of the transformer encoder to the thermal conductivity of COFs. We used the mean squared error (MSE) as a loss function for training. The dataset of 2,471 COFs is randomly split into two subsets: 90% of the data is used for training (out of which 10% for validation), and the remaining data is used for testing. Additional details of the training protocol are provided in Supporting Information Section 3.
We demonstrate the model’s prediction accuracy on the test dataset in Figure 3, where it achieves a goodness-of-fit $R^{2}$ of $0.909\pm 0.006$ (across five random seeds) and a mean absolute error (MAE) of $0.075\ \mathrm{Wm^{-1}K^{-1}}$ for predicting $\kappa$. The model’s strong performance suggest that there may be gaps in the feature sets typically used to describe COFs. It also highlights the possibility that additional key descriptors, beyond those commonly used, could play an important role in predicting thermal conductivity, thus warranting further exploration of structure-property relationships.
Discovering novel thermal transfer mechanisms via self-attention
We leverage the multi-head self-attention mechanism [50] of the transformer architecture to interpret the predictions made by the deep learning model. The attention mechanism in transformers is designed to identify and assign weights, known as attention scores, to the significance of different parts of the input data/encodings relative to all the other parts. When applied to molecular structures, such as COFs, the attention mechanism enables the model to dynamically focus attention on key molecular substructures in relation to all the other substructures present based on their relevance to the property being predicted, in this case, thermal conductivity. In multi-head attention, multiple attention heads are needed to capture different aspects or relationships within the data simultaneously. Specifically, in the context of the PMTransformer model applied to a COF to predict its thermal conductivity, we calculate the attention for each atom in the COF by averaging the attention scores from all attention heads and computing the joint attention via multiplicative aggregation across all layers of the transformer encoder.
Additionally, we identify the main branch of a COF as the shortest continuous path connecting the boundary points necessary for periodicity. We then classify all the atoms extending and excluded from the main branch as dangling mass. More details on how we classify each atom as either part of the main branch or part of a dangling side branch are presented in Supporting Information Section 4. We observe distinct patterns in the attention assigned to various atomic sites within a COF’s graph representation, which correspond to the location of the dangling masses. Two representative examples are shown in Figure 4.
Consider the first example (see Figure 4a-b). The top row illustrates a COF structure (Figure 4a) from the test dataset with a thermal conductivity of $1.105\ \mathrm{Wm^{-1}K^{-1}}$ and no dangling masses except hydrogen atoms. Additionally, the atom-wise attention scores are uniformly distributed and low for all the carbon atoms with some elevated attention to the nitrogen atoms. In the second row, we illustrate a COF structure (Figure 4b) with the same topology as the previous one, but with a much lower thermal conductivity of $0.533\ \mathrm{Wm^{-1}K^{-1}}$. Its attention profile reveals that certain atom groups are being paid special attention to by the transformer, i.e., specifically the $-\text{NO}_{2}$ functional group on the benzene ring that is dangling from the main branch of the COF. Notably, the same groups of atoms that exhibit higher attention scores are also classified as dangling masses (see Figure 4b attention profile & dangling mass).
@@
The same trend is observed for numerous examples, with additional ones presented in the Supporting Information Section 5. We conclude that this increased attention is indicative of the deep learning model’s understanding that these dangling masses are significant in predicting thermal conductivity. Furthermore, it suggests that the presence of dangling masses disrupts the heat transfer pathways and thereby reduces the thermal conductivity through the material. A similar effect of dangling mass lowering the thermal conductivity has been reported previously in singular polymer chains [54] and amorphous polymers [55], but remains a novel predictor for COFs which are ordered polymeric structures.
@@
To further validate the structure-property relationship interpreted from the deep learning model, we examine the impact of dangling atoms on the vibrational density of states (VDOS) for the aforementioned contrasting COF examples. VDOS, which characterizes the distribution of vibrational modes in a system as a function of frequency, is known to impact thermal transfer properties. Overlaps in VDOS profiles between different atoms have been shown to affect these properties in both COFs [56] and MOFs [57].
Figure 4: Two example pairs (contained in dashed boxes) of COFs with same topologies, similar geometric descriptors, but contrasting thermal conductivities. The first column illustrates the COF structures. The second column shows the atom-level attention score profile computed by the attention mechanism. The third column shows the same COF structure distinguishing atoms on the main branch and dangling atoms (with separate distinction for hydrogen atoms). The fourth column shows the VDOS profiles of various groups of atoms within the corresponding COF structure with the overlap metric $S$. The legend indicates the VDOS profile for main branch atoms (.)(.)( . ) and dangling atoms (.$(.)^{\text{(d)}}$.
@@
Through MD simulations, the VDOS is calculated using the Fourier transform of the normalized velocity autocorrelation function of specific groups of atoms (see Supporting Information Section 6 for details). Figure 4 (third column) show the VDOS profile of carbon, nitrogen, and oxygen atoms in the representative COF examples. For any COF, we define a VDOS overlap metric $S$ as the ratio of the area under the curve (AUC) of the minimum VDOS across all atom types (at each frequency) and the AUC of the maximum VDOS across all atom types (at each frequency), i.e.,
$S=\frac{\int_{\omega}\min\left\{\ \mathrm{f^{V}_{C}(\omega),\ f^{V}_{N}(\omega),\ f^{V}_{O}(\omega),\ f^{V}_{B}(\omega),\ f^{V,d}_{C}(\omega),\ f^{V,d}_{N}(\omega),\ f^{V,d}_{O}(\omega),\ f^{V,d}_{B}(\omega)\ }\right\}\;\!\mathrm{d}\omega}{\int_{\omega}\max\left\{\ \mathrm{f^{V}_{C}(\omega),\ f^{V}_{N}(\omega),\ f^{V}_{O}(\omega),\ f^{V}_{B}(\omega),\ f^{V,d}_{C}(\omega),\ f^{V,d}_{N}(\omega),\ f^{V,d}_{O}(\omega),\ f^{V,d}_{B}(\omega)}\right\}\;\!\mathrm{d}\omega}.$ (1)
@@
Here, $\omega$ denotes the frequency; $\mathrm{f_{C}^{V}}(\omega)$, $\mathrm{f_{N}^{V}}(\omega)$, $\mathrm{f_{O}^{V}}(\omega)$, and $\mathrm{f_{B}^{V}}(\omega)$ denote the VDOS at frequency $\omega$ for the carbon, nitrogen, oxygen and boron atoms, respectively. Additionally, we differentiate between atoms in the main branch and dangling atoms, where the VDOS profiles for dangling atoms are denoted with a (.$(.)^{\text{d}}$ superscript. A large overlap in VDOS profiles of different atom types would result in $S\approx 1$, whereas a small overlap would yield $S\approx 0$. A high overlap indicates that phonon waves (i.e., quantized modes of vibrations responsible for thermal energy transfer within the crystal lattice) can propagate freely throughout the structure. This would lead to minimized phonon scattering and facilitate efficient heat transfer across the material [58].
@@
In the COF structures, which lack or have minimal dangling atoms, the VDOS profiles of the atoms are broad and overlap significantly, e.g., $S\approx 0.41$ and $0.49$ for examples in Figure 4a and c, respectively. This results in an even distribution of vibrational modes across a wide frequency spectrum for all atoms. In contrast, the COFs with substantial dangling atoms exhibit a VDOS profile with minimal overlap, e.g., $S\approx 0.08$ and $0.09$ for examples in Figure 4b and d, respectively. The vibrational modes of individual atoms in these COFs are confined to much narrower frequency bands, which leads to a less harmonized VDOS profile. The lack of overlap suggests that the vibrations are highly localized around the dangling atoms, which causes the phonon waves to be scattered and disrupting their ability to transfer heat efficiently through the material [59]. We interpret the mismatch in the VDOS profiles as an energy barrier for phonon transport, where phonons encounter resistance, preventing smooth energy transfer across atoms. This suggests that the absence of dangling mass is crucial for enhancing thermal conductivity in COFs. Dangling atoms create these mismatches in VDOS, which in turn hinders phonon transport and reduces thermal conductivity.
@@
To further confirm the effect of dangling atoms on phonon dispersion in COFs, we calculate the phonon spectral energy density (pSED) of the example COFs and their dangling counterparts, as presented in Section 5 in Supporting Information. The pSED of COFs containing more dangling atoms exhibits higher magnitudes and broadening bands, indicating stronger anharmonicity and vibrational scattering in the vibrational modes, consequently resulting in reduced thermal conductivity [60].
@@
where $m_{i}$ is the mass of the $i^{\text{th}}$ atom belonging to the set of either dangling atoms or of non-dangling main branch atoms – denoted by $\mathcal{D}$ and $\mathcal{M}$, respectively. Higher DMR values indicate a greater proportion of dangling masses relative to the COF’s total mass.
We then introduce DMR as an additional feature, alongside pore size, density, void fraction, and surface area, and train classical ensemble regression models, such as the Random Forest [61], Gradient Boosting [62], XGBoost [63], and AdaBoost [64] and outline their performances in Table 2. To assess the importance of individual features, including DMR, we analyzed their Gini importance and conducted a permutation feature importance analysis [65]. These values quantify the contribution of each feature to the overall predictive power of the model. The Gini importance measures how much each feature reduces the Gini impurity or randomness when making predictions in the ensemble model [61]. On the other hand, permutation feature importance measures the impact of a feature by assessing how much the model’s performance decreases when the values of that feature are randomly shuffled. A greater drop in performance indicates that the feature is more important in predicting the target variable [61]. We note that for the best performing regressor, the Random Forest, DMR emerges as the second most influential predictor with 27.8% Gini importance, closely following behind density (see Table 2). Similarly, the permutation feature importance analysis also highlights DMR as a key feature, further supporting its significant role in predicting thermal conductivity, with its importance consistently ranking just below density. Finally, we conducted SHAP analysis [66] on the the Random Forest to better understand the impact of each feature on the individual test data points. As shown in Figure 5, the analysis reveals that higher DMR values are associated with lower predicted $\kappa$, confirming the trend we observed. The consistency of these findings across different methods—Gini, permutation feature importance, and SHAP—reinforces the importance of DMR as a key factor in predicting thermal conductivity.
@@
Figure 5: Distribution of the effect of each feature on the model’s predictions. Each point represents a SHAP value for a feature across all samples, with the color intensity indicating feature values (ranging between minimum and maximum values). The position along the x-axis represents the magnitude of the feature’s impact.
@@
Generalizability and high-throughput screening for discovering COFs with tailored properties
@@
Figure 6: High-throughput screening. (a) Distribution of $\kappa$ in the training dataset. Dashed lines indicated minimum, mean, and maximum $\kappa$ across the training dataset. (b,c) Distribution of $\kappa$ of selected COF candidates identified in high-throughput screening for low and high $\kappa$ in the unlabeled datasets of (b) Mercado et al. [47] and (c) Lan et al. [67] While the former dataset shares the same design space as the training dataset, the latter has a different design space. All thermal conductivities shown here are computed using MD. (d) Representative COFs identified through high-throughput screening with $\kappa$ lower and higher than the minimum and maximum, respectively, across the training dataset. Also indicated are their corresponding DMR and $S$ values.
@@
To evaluate the generalization capabilities of the deep learning model for high-throughput screening, we screen previously unseen 6,170 two-dimensional COFs in the unlabelled dataset by Mercado et al. [47] from which our training and test sets were originally subsampled. Moreover, we screen $35,638$ two-dimensional COFs in the unlabelled dataset by Lan et al. [67] This dataset covers a vastly different design space with a variety of linkers and knots previously unseen by our trained model.
@@
The screening is performed with the objective of facilitating the identification COFs with extremely high or low thermal conductivity without resorting to computationally expensive simulations. We observe an average screening rate of 0.07 seconds per COF structure (see Supporting Information Section 7 for details on computational runtimes), thereby enabling a speedup of almost seven orders of magnitude compared to MD-based screening.
@@
Once potential candidates with extreme thermal conductivities are identified using the PMTransformer, we then calculate their thermal conductivities through MD simulations for these selected few, rather than for the entire datasets. The results are summarized in Figure 6. Our screening process on the dataset by Mercado et al. [47] revealed COFs exhibiting thermal conductivities up to twice the maximum value observed in the training set, but stemming from the same design space. By screening the dataset by Lan et al. [67], we discover COFs with thermal conductivities five times as high as the ones in our training set. Additionally, we discovered several COFs with thermal conductivities lower than any values previously recorded in our dataset. A visual inspection reveals that these COFs feature a combination of large pore sizes, low density, and a substantial presence of dangling masses. For all the identified COFs, we observe that thermal conductivity shows a strong trend with both dangling mass ratio (DMR) and VDOS overlap ratio ($S$), which is in agreement with the previous observations.

### v3

@@ Results
In this study, we leveraged the pre-trained transformer encoder of the PMTransformer model to perform transfer learning for the prediction of the thermal conductivity of COFs. We fine-tune the transformer model using the pre-trained weights as a starting point and jointly train a prediction head based on a multi-layer perceptron architecture to map the output of the transformer encoder to the thermal conductivity of COFs. We used the mean squared error (MSE) as a loss function for training. The dataset of 2,471 COFs is randomly split into two subsets: 90% of the data is used for training (out of which 10% for validation), and the remaining data is used for testing. We repeat the training across five different random seeds, where each seed initializes the prediction head randomly while keeping the transformer initialized with the same pre-trained weights. Additional details of the training protocol and an ablation study are provided in Supplementary Information Section 3.
@@
Discovering novel thermal transfer mechanisms for two-dimensional COFs via self-attention
We leverage the multi-head self-attention mechanism 49 of the transformer architecture to interpret the predictions made by the deep learning model. The attention mechanism in transformers is designed to identify and assign weights, known as attention scores, to the significance of different parts of the input data/encodings relative to all the other parts. When applied to molecular structures, such as COFs, the attention mechanism enables the model to dynamically focus attention on key molecular substructures in relation to all the other substructures present based on their relevance to the property being predicted, in this case, thermal conductivity. In multi-head attention, multiple attention heads are needed to capture different aspects or relationships within the data simultaneously. Specifically, in the context of the PMTransformer model applied to a COF to predict its thermal conductivity, we calculate the attention for each atom in the COF by averaging the attention scores from all attention heads and computing the joint attention via multiplicative aggregation across all layers of the transformer encoder.
Additionally, we identify the main branch of a COF as the shortest continuous path connecting the boundary points necessary for periodicity. We then classify all the atoms extending and excluded from the main branch as dangling mass. More details on how we classify each atom as either part of the main branch or part of a dangling side branch are presented in Supplementary Information Section 4. We observe distinct patterns in the attention assigned to various atomic sites within a COF’s graph representation, which correspond to the location of the dangling masses. Two representative examples are shown in Fig. 4 and Fig. 5.
[…]
@@
The same trend is observed for numerous examples, with additional ones presented in the Supplementary Information Section 5. We hypothesize that this increased attention is indicative of the deep learning model’s understanding that these dangling masses are significant in predicting thermal conductivity. Furthermore, it suggests that the presence of dangling masses disrupts the heat transfer pathways and thereby reduces the thermal conductivity through the material. A similar effect of dangling mass lowering the thermal conductivity has been reported previously in singular polymer chains 53 and amorphous polymers 54. This effect, while known in disordered systems, has not been examined in crystalline COFs. Here, we demonstrate that it remains relevant even in ordered, porous frameworks, inviting future investigation with high-fidelity modeling methods (such as density functional theory) to examine the underlying mechanism in more detail.
@@
To further validate the structure-property relationship interpreted from the deep learning model, we examine the impact of dangling atoms on the vibrational density of states (VDOS) for the aforementioned contrasting COF examples. VDOS, which characterizes the distribution of vibrational modes in a system as a function of frequency, is known to impact thermal transfer properties. Overlaps in VDOS profiles between different atoms have been shown to affect these properties in both COFs 55 and MOFs 56.
@@
Through MD simulations, the VDOS is calculated using the Fourier transform of the normalized velocity autocorrelation function of specific groups of atoms (see Supplementary Information Section 6 for details). Fig. 4 and Fig. 5(third column) show the VDOS profile of carbon, nitrogen, and oxygen atoms in the representative COF examples. For any COF, we define a VDOS overlap metric $S$ as the ratio of the area under the curve (AUC) of the minimum VDOS across all atom types (at each frequency) and the AUC of the maximum VDOS across all atom types (at each frequency), i.e.,
$\begin{split}S=\frac{\int_{\omega}\min\left\{\ \mathrm{f^{V}_{C}(\omega),\ f^{V}_{N}(\omega),\ f^{V}_{O}(\omega),\ f^{V}_{B}(\omega),\ f^{V,d}_{C}(\omega),\ f^{V,d}_{N}(\omega),\ f^{V,d}_{O}(\omega),\ f^{V,d}_{B}(\omega)\ }\right\}\;\!\mathrm{d}\omega}{\int_{\omega}\max\left\{\ \mathrm{f^{V}_{C}(\omega),\ f^{V}_{N}(\omega),\ f^{V}_{O}(\omega),\ f^{V}_{B}(\omega),\ f^{V,d}_{C}(\omega),\ f^{V,d}_{N}(\omega),\ f^{V,d}_{O}(\omega),\ f^{V,d}_{B}(\omega)}\right\}\;\!\mathrm{d}\omega}.\end{split}$ (1)
@@
Here, $\omega$ denotes the frequency; $\mathrm{f_{C}^{V}}(\omega)$, $\mathrm{f_{N}^{V}}(\omega)$, $\mathrm{f_{O}^{V}}(\omega)$, and $\mathrm{f_{B}^{V}}(\omega)$ denote the VDOS at frequency $\omega$ for the carbon, nitrogen, oxygen and boron atoms, respectively. Additionally, we differentiate between atoms in the main branch and dangling atoms, where the VDOS profiles for dangling atoms are denoted with a $(.)^{\text{d}}$ superscript. A large overlap in VDOS profiles of different atom types would result in $S\approx 1$, whereas a small overlap would yield $S\approx 0$. A high overlap indicates that phonon waves (i.e., quantized modes of vibrations responsible for thermal energy transfer within the crystal lattice) can propagate freely throughout the structure. This would lead to minimized phonon scattering and facilitate efficient heat transfer across the material 57.
@@
In the COF structures, which lack or have minimal dangling atoms, the VDOS profiles of the atoms are broad and overlap significantly, e.g., $S\approx 0.41$ and $0.49$ for examples in Fig. 4a and Fig. 5a, respectively. This results in an even distribution of vibrational modes across a wide frequency spectrum for all atoms. In contrast, the COFs with substantial dangling atoms exhibit a VDOS profile with minimal overlap, e.g., $S\approx 0.08$ and $0.09$ for examples in Fig. 4b and Fig. 5b, respectively. Notably, dangling atoms significantly reduce the VDOS overlap in both low-frequency (0 to 20 THz) and high-frequency (40 to 60 THz) modes. The vibrational modes of individual atoms in these COFs are confined to much narrower frequency bands, which leads to a less harmonized VDOS profile. The lack of overlap suggests that the vibrations are highly localized around the dangling atoms, which causes the phonon waves to be scattered and disrupting their ability to transfer heat efficiently through the material 58. We interpret the mismatch in the VDOS profiles as an energy barrier for phonon transport, where phonons encounter resistance, preventing smooth energy transfer across atoms. This suggests that the absence of dangling mass is crucial for enhancing thermal conductivity in COFs. Dangling atoms create these mismatches in VDOS, which in turn hinders phonon transport and reduces thermal conductivity. A similar trend is observed in additional pairs of COFs (Fig. S9, Supplementary Information) and the correlation between $\kappa$ and $S$ for a selected representative set of eight COFs is presented in Fig. S6c (Supplementary Information).
Fig. 4: An example pair of COFs with same topologies, similar geometric descriptors, but contrasting thermal conductivities. The first column illustrates the COF structures. The second column shows the atom-level attention score profile computed by the attention mechanism. The third column shows the same COF structure distinguishing atoms on the main branch and dangling atoms (with separate distinction for hydrogen atoms). The fourth column shows the VDOS profiles of various groups of atoms within the corresponding COF structure with the overlap metric $S$. The legend indicates the VDOS profile for main branch atoms $(.)$ and dangling atoms $(.)^{\text{(d)}}$. The reported thermal conductivities are obtained from NEMD simulations, rather than predicted by the PMTransformer model.
Fig. 5: An example pair of COFs with same topologies, similar geometric descriptors, but contrasting thermal conductivities. The first column illustrates the COF structures. The second column shows the atom-level attention score profile computed by the attention mechanism. The third column shows the same COF structure distinguishing atoms on the main branch and dangling atoms (with separate distinction for hydrogen atoms). The fourth column shows the VDOS profiles of various groups of atoms within the corresponding COF structure with the overlap metric $S$. The legend indicates the VDOS profile for main branch atoms $(.)$ and dangling atoms $(.)^{\text{(d)}}$. The reported thermal conductivities are obtained from NEMD simulations, rather than predicted by the PMTransformer model.
@@
To further confirm the effect of dangling atoms on phonon dispersion in COFs, we calculate the phonon spectral energy density (pSED) of the example COFs and their dangling counterparts, as presented in Section 5 in Supplementary Information. The pSED of COFs containing more dangling atoms exhibits higher magnitudes and broadening bands, indicating stronger anharmonicity and vibrational scattering in the vibrational modes, consequently resulting in reduced thermal conductivity.
@@
where $m_{i}$ is the mass of the $i^{\text{th}}$ atom belonging to the set of either dangling atoms or of non-dangling main branch atoms – denoted by $\mathcal{D}$ and $\mathcal{M}$, respectively. Higher DMR values indicate a greater proportion of dangling masses relative to the COF’s total mass. In Supplementary Information Section 1.3, we qualitatively show the relationship between DMR, the VDOS overlap $S$, and $\kappa$.
We then introduce DMR as an additional feature, alongside pore size, density, void fraction, and surface area, and train classical ensemble regression models, such as the Random Forest 59, Gradient Boosting 60, XGBoost 61, and AdaBoost 62 and outline their performances in Table 2. These simpler models are employed not to compete with the PMTransformer in predictive accuracy, but to provide directly interpretable feature importance scores and understanding the relevance of DMR. In contrast, extracting meaningful and quantitative feature importance from the PMTransformer is non-trivial due to its complex model architecture and the absence of an explicit correspondence between input features and learned representations. To assess the importance of individual features, including DMR, we analyzed their Gini importance and conducted a permutation feature importance analysis 63. These values quantify the contribution of each feature to the overall predictive power of the model. The Gini importance measures how much each feature reduces the Gini impurity or randomness when making predictions in the ensemble model 59. On the other hand, permutation feature importance measures the impact of a feature by assessing how much the model’s performance decreases when the values of that feature are randomly shuffled. A greater drop in performance indicates that the feature is more important in predicting the target variable 59. We note that for the best performing regressor, the Random Forest, DMR emerges as the second most influential predictor with 27.8% Gini importance, closely following behind density (see Table 2). Similarly, the permutation feature importance analysis also highlights DMR as a key feature, further supporting its significant role in predicting thermal conductivity, with its importance consistently ranking just below density. Finally, we conducted SHAP analysis 64 on the the Random Forest to better understand the impact of each feature on the individual test data points. As shown in Fig. 6, the analysis reveals that higher DMR values are associated with lower predicted $\kappa$, confirming the trend we observed. The consistency of these findings across different methods—Gini, permutation feature importance, and SHAP—reinforces the importance of DMR as a key factor in predicting thermal conductivity. We note that the computed importance scores are inherently linked to the design space and dataset introduced by Mercado et al. 47. Specifically, if the dataset did not contain structures with dangling masses, the DMR feature would naturally show little to no importance. Thus, the relevance of DMR as a descriptor depends on its variability within the dataset under consideration.
@@
Fig. 6: Distribution of the effect of each feature on the model’s predictions. Each point represents a SHAP value for a feature across all samples, with the color intensity indicating feature values (ranging between minimum and maximum values). The position along the x-axis represents the magnitude of the feature’s impact.
@@
Generalizability and high-throughput screening for discovering two-dimensional monolayer COFs with tailored properties
@@
Fig. 7: High-throughput screening. (a) Distribution of $\kappa$ in the training dataset. Dashed lines indicated minimum, mean, and maximum $\kappa$ across the training dataset. (b,c) Distribution of $\kappa$ of selected COF candidates identified in high-throughput screening for low and high $\kappa$ in the unlabeled datasets of (b) Mercado et al. 47 and (c) Lan et al. 65 While the former dataset shares the same design space as the training dataset, the latter has a different design space. All thermal conductivities shown here are computed using MD. (d) Representative COFs identified through high-throughput screening with $\kappa$ lower and higher than the minimum and maximum, respectively, across the training dataset. Also indicated are their corresponding DMR and $S$ values.
@@
To evaluate the generalization capabilities of the deep learning model for high-throughput screening, we screen previously unseen 6,170 two-dimensional COFs in the unlabelled dataset by Mercado et al. 47 from which our training and test sets were originally subsampled. Moreover, we screen $35,638$ two-dimensional COFs in the unlabelled dataset by Lan et al. 65 This dataset covers a vastly different design space with a variety of linkers and knots previously unseen by our trained model.
@@
The screening is performed with the objective of facilitating the identification COFs with extremely high or low thermal conductivity without resorting to computationally expensive simulations. We observe an average screening rate of 0.07 seconds per COF structure (see Supplementary Information Section 7 for details on computational runtimes), thereby enabling a speedup of almost seven orders of magnitude compared to MD-based screening.
@@
Once potential candidates with extreme thermal conductivities are identified using the PMTransformer, we then calculate their thermal conductivities through MD simulations for these selected few, rather than for the entire datasets. This approach circumvents reliance on the PMTransformer’s absolute predictions by validating promising candidates with high-fidelity MD calculations. The results are summarized in Fig. 7. Our screening process on the dataset by Mercado et al. 47 revealed COFs exhibiting thermal conductivities up to twice the maximum value observed in the training set, but stemming from the same design space. By screening the dataset by Lan et al. 65, we discover COFs with thermal conductivities five times as high as the ones in our training set. Additionally, we discovered several COFs with thermal conductivities lower than any values previously recorded in our dataset. A visual inspection reveals that these COFs feature a combination of large pore sizes, low density, and a substantial presence of dangling masses. For all the identified COFs, we observe that thermal conductivity shows a strong trend with both dangling mass ratio (DMR) and VDOS overlap ratio ($S$), which is in agreement with the previous observations.